OpenStudy (anonymous):

Factor completely 49x2 – 81 ?

OpenStudy (anonymous):

Use \(a^2-b^2=(a+b)(a-b).\)

OpenStudy (anonymous):

What..?

OpenStudy (anonymous):

Let \(a^2=49x^2\) and \(b^2=81.\) Now what is \((a+b)(a-b)?\)

OpenStudy (anonymous):

How do I plug it in?